by TEI TEI

04/10/08 – Triton submitted closeout paperwork on this project. TTC review and approved. TEi has to write a change order deduct for the allowance money not used on this project and will send that to Triton for processing on this project.

12/28/07 – Grady asked about payapp # 2 and TEI reported that punchlist, O&M’s, payapp 1 have all been received and sent. To closeout this project we need closeout documents, final payapp, certified payroll and an as-built dwg.

12/21/07 – O&M’s have been received by TEI. TTC approved them for distribution to the County. We are sending them to Bert Watts. TTC asked for an additional airflow balancing on the main units themselves for maximum flow with the VFD at 100% to included later.

12/19/07 – JS reported today that the final punchlist work is complete. TEI can now backcheck the punchlist and closeout the job. MB asked about allowance money remaining for moving a piece of equipment at the CH.

12/12/07 – John Sanders faxed over a final signed punchlist to TTC for review yesterday and reported today that the flow switch cut outs have been approved by the code officials. They will now get them installed into the units for final inspection and approval by the city.

11/30/07 – Triton and Honeywell have finished all their punchlist items but the project did not pass final inspection because of the new mechanical code requirement about secondary condensate drainage under cooling coils. This requirement is listed under the General Regulations Section 307 of the OMC for 2007. TTC sent scanned copies of the code to Mike B and Grady R for assistance in getting the permit and finishing this job. Triton reports, “We are getting documentation from carrier as to what they recommend as an overflow protection device to shutdown in the event there is water present outside of the drain pan. The inspector is ok with this as long as carrier is ok with the install – Grady Reed.

10/17/07 – John Sanders called today to tell us that Triton is lowering the piping on AHU-360 and will coordinate that with MB, bringing out a sheet metal cover for the insulation touching the door on CR 360, moving the CHWS label to the correct line and that they have finished the painting of the doors which I told him was fine they way they looked now, . The controls are still being finished

10/08/07 – TTC sent a revised punchlist with picture email to Grady based on his conversation with Mike Bartlett and Willie Schadler. Triton is trying to finish this project but balancing and controls remain to be done.

10/0407 – County said AHU piping was not switched on original air handler. They may mean the piping blocking the filter rack on the second air handler. Regardless, Triton has to change the piping to match the design.

10/01/07 – Triton has stated that have finished punchlist work and controls are finishing their. Payapp # 1 has been received. Grady has the two VFD’s that TEI is supposed to purchase for our error on this job. Units are installed and working very well as far as I know.

09/12/07 – Triton and QCE are on site performing punchlist items. Honeywell has final control parts to install so system can begin running the VFD’s automatically. Randy from Honeywell said he should be able to install the components this week. TTC will check back on progress next week. The CHW piping has not been relocated yet.09/10/07 – TTC inquired about the status of this project with contractors today.

09/03/07 – TTC finished and sent the punchlist to Triton for this project. Controls are not finished and VFD’s are still in manual causing some overcooling of the space. I am confident that the systems will work when complete.

08/22/07 – Both units are now installed in the space. There were some problems on the initial installation but they have now been worked out. The VFD’s were ordered at 480V when the building really has 208V at the units. Triton quick shipped 2 new units that week and the 340 unit was running the end of the week. Air noise and unit noise has been drastically reduced with the new unit but the motor was making a hum that was fixed by changing the tap voltage and reprogramming the VFD. The County asked for new isolation valves in back of the old ones and Triton has provided a change order cost for this. TTC forwarded to County and is awaiting a decision on allowance usage. The CHWS/R piping was connected backwards to the 340 units but Triton flipped it during the 360 changeout work. The 360 changeout work went much better but MB reported that the new CHW piping interferes with filter rack removal. Grady said that Triton will take care of this. MB reported that the motor noise was back and that the judges both complained. Grady had the VFD rep make some changes and thought it was fine. TTC will contact MB to check. TEI is waiting until the controls are finished before punchlisting the job which should be next week. We have coordinated this with MB. TTC told Triton that the dwgs do NOT specifically call out for the steam trap, isolation valve and strainer to be replaced. It was a lineweight issue with Queen City printing that showed them dark.

07/31/07 – TTC recommended the 340 AHU piping be changed during the 360 shutdown weekend work. We can coordinate it so that we do 360 AHU first and if there is time left we can do the 340 AHU. QCE will also change the 30A fuses to 25A during the shutdown. Carrier will investigate the harmonics of the motor noise to get that removed from the unit. AHU seems to be cooling the courtroom just fine.

07/30/07 – NR verified the fuse and wire sizing for QCE. There only modification from the original design is that the disconnect switch frame is larger than intended. The fuse and wire size match the design so the frame size is acceptable.

07/27/07 – After a week of tribulations, the 340 AHU is up and running. There are several issues that have happened on this project that have this changeout tough. TEI spec’d the wrong voltage on the VFD and had the CHWS and CHWR labeled backwards on the unit. Triton did not have all the material they needed. The unit is whisper quiet but the motor makes a funny noise that I think is not right. We are working on resolving these issues and will get this unit checked and balanced before we changeout the 360 unit.

07/23/07 – Triton informed everyone today that TEI spec’d 480V VFD’s while the correct supply voltage to the units is 230V. TTC checked the drawings and we do call for 480V. We don’t know how this happened because we checked this several times during design and coordinated amongst ourselves. Regardless, we have to find 230V VFD’s from Carrier immediately to keep the project moving forward. The old 340 air handler has been removed and Triton was to remove it from site yesterday. The new unit is installed but not powered up yet. Triton is working on the new drive as we speak.

07/18/07 – Triton submitted a $2300 allowance charge for replacing all pneumatic actuators with electric actuators. AKM approved and TTC sent out the paperwork and approved the change. We are scheduled to have a coordination meeting on Friday with Triton for last minute changeout details for the new unit.

06/25/07 – We had our first construction mtg yesterday with our new project manager Grady Reed. John Sanders is going to be our field foreman on this job. We walked the site and Triton is now putting together a schedule and employee list. Mike is checking on best weekends for the judges. TTC is going to distribute the meeting notes for the project.

05/30/07 – Equipment is scheduled to be delivered in mid-June for immediate installation into project.

05/17/05 – TEI finished the control submittals and sent them back to Grady Reed of Triton Services.

03/15/07 – Submittals have been returned to Triton for ordering on this project. Controls asked about 120V power and NR answered via email that 4th floor is fine.

02/21/07 – Pre Const Mtg yesterday. Meeting notes today. PDF

09/01/06 – TTC met Tim Centers on site today to discuss fitting the unit into the space. The only door from the hallway has a maximum width of 35”. The unit must be completely broke down and re-assembled in place. This is possible with the Carrier 39M series although the entire unit must be taken apart. Habegger reported that the unit will be quieter especially since the fan and all associated components are located inside the unit. We are buying a VFD with the unit for speed control. Also we discovered today that the unit when constructed is tall enough so the fan discharge will match up with the existing supply ductwork. There is no need to put a supply elbow on top of the unit. We are also going to keep the existing return and OA as is and butt our unit up against the existing filter rack. There is just enough room for the transition to fit. TTC went over the controls with GD today. The chilled water valve and actuator are new. The steam valve and its actuator may be replaced. The OA damper will remain in use as is.